OpenStudy (rainbow_rocks03):

PlZ help me to understand how to get my answer. Evaluate. (16 + 32 ÷ 4)(21 – 6 • 3)

OpenStudy (rainbow_rocks03):

No direct answers

OpenStudy (rainbow_rocks03):

@horsegal244

OpenStudy (horsegal244):

Just Answer it how it says to but do this first (16 + 32 ÷ 4)

OpenStudy (rainbow_rocks03):

16+32=48/4=12

OpenStudy (horsegal244):

Yes now do (21 – 6 • 3)

OpenStudy (rainbow_rocks03):

21-6=15*3=45

OpenStudy (horsegal244):

ok now put the 2 together

OpenStudy (rainbow_rocks03):

add or multiply?

OpenStudy (vivi_):

72

OpenStudy (rainbow_rocks03):

how?

OpenStudy (vivi_):

Use Mathway. Its helps a lot. Its never wrong

OpenStudy (horsegal244):

Add

OpenStudy (rainbow_rocks03):

45+12=57

OpenStudy (horsegal244):

Yep Not every website is right but you are right

OpenStudy (rainbow_rocks03):

@horsegal244 TY for your help , but that was what I got and it still marked it wrong. It says the answer is 72. I don't understand how.

OpenStudy (horsegal244):

That is strange let me do the math once more

OpenStudy (rainbow_rocks03):

ok TY

OpenStudy (horsegal244):

oh i totally forgot that it was multiplying sorry...

OpenStudy (rainbow_rocks03):

oh!

OpenStudy (rainbow_rocks03):

NP TY so much

OpenStudy (horsegal244):

Welcome

OpenStudy (rainbow_rocks03):

where did you multiply?

OpenStudy (vivi_):

Because it P.E.M.D.A.S (16 + 32 / 4) (21 - 6 * 3) (16 +32 / 4) (21 - 18) (24) 3 then you multiply 24 and 3 and you get 72